OpenStudy (anonymous):

Simplify sqrt75-4sqrt8-3sqrt32

OpenStudy (anonymous):

sqrt 75 = sqrt (25*3) = 5sqrt 3 4sqrt 8 = 4 sqrt (4*2) = 8 sqrt 2 3sqrt 32 = 3 sqrt (16*2) = 12sqrt2

OpenStudy (anonymous):